The following line was given in a solution to a question.
P(A)= (6C3 x 43C3 ) + ( 6C4 x 43C2 ) divided by 49 6

What might the question have been? Be sure to identify and explain everything that you know is true about the question.
1
Expert's answer
2020-07-13T19:01:54-0400

Solution:

Take for example Canadian lottery 6/49.


The question could be as follow:

"Let A be event that there are 3 winning numbers and 3 losing numbers or

4 winning numbers and 2 losing numbers. Determine the probability of event A"


Then P(A)=m/n.


6C3 - number of ways to draw 3 numbers from 6 winning numbers,

43C3 - number of ways to draw 3 numbers from 43 losing numbers,

6C4 - number of ways to draw 4 numbers from 6 winning numbers,

43C2 - number of ways to draw 2 numbers from 43 losing numbers,

m=6C3 x 43C3 +6C4 x 43C2.


49C6 - number of ways to draw 6 numbers from 49 numbers,

n=49C6


P(A)=(6C3 x 43C3 +6C4 x 43C2)/49C6=

=(6!/3!/3!*43!/3!/40!+6!/4!/2!*43!/2!/41!)/(49!/6!/43!)=37195/1997688.


Here nCk=n!/(k!(n-k)!) is binomial coefficient.